Find a particular solution of the given differential equation. Use a CAS as an aid in carrying out differentiations, simplifications, and algebra.
y(4) + 2y'' + y = 6 cos x − 8x sin x

yp =
1
Expert's answer
2020-08-03T18:41:43-0400

"\lambda^4+2\cdot \lambda^2+1=0\\
(\lambda^2+1)^2=(\lambda+i)^2\cdot (\lambda-i)^2=0\\
\lambda_1=\lambda_2=-i, \lambda_3=\lambda_4=i\\
y_{h}=(c_1+c_2x)sin(x)+(c_3+c_4x)cos(x)\\
y_0=(\frac{1}{4}x^2-\frac{9}{16})cos(x)+(\frac{1}{3}x^3-\frac{1}{2}x)sin(x).\\
y(x)=y_{h}+y_{0}=(c_1+c_2x)sin(x)+(c_3+c_4x)cos(x)+
(\frac{1}{4}x^2-\frac{9}{16})cos(x)+(\frac{1}{3}x^3-\frac{1}{2}x)sin(x);"